OpenStudy (anonymous):

Can someone please explain how to find what percent of 42 is 28? Just explanation...thanks!

hero (hero):

x percent of 42 = 28 is/of = %/100 28/42 = x/100 Cross multiply: 2800 = 42x 2800/42 = x

OpenStudy (anonymous):

Thanks!!!